OpenStudy (anonymous):

I really am confused on this problem and have no idea how to even do it. Can anyone please help? f(x)=x+10/x-9 1) is point 4, - 15/4 function on graph f 2) if f=x-2 what is f(x) 3) list point on graph of f 4) if f(x)=2, what is x? 5) Domain 6) any intercepts

OpenStudy (anonymous):

remember that f(x) = y so for each value of x that your taking or getting a value of "y" or "image" as we call it here... so I think the first step should be finding your domain of x, that is where x excists on your function notice that u have the denominator x-9 and remember that you cannot divide anything by 0, so you have to say your domain is: <Domain for f(x)=x+10/x-9 / all reals minus X-9=0 > or just say x cannot be equal to 9 will keep writting read this for the moment ..

OpenStudy (anonymous):

now for number 1 you have to see if what they're saying is right. they are saying that (x,y) = (4,- 15/4) is a point on the function here you have to see if that's right remember you can writte ur function as y=x+10/x-9 here you can either replace the x or the y for the numbers given on the question, it doesn't really mather, what mathers is that you get the same number for the other variable, so i'll go with replacing the x, and I should get y=- 15/4. If you replace the y you should get the value for x = 4. if not the case then the point doesn`t exist on our graph... y=(x+10)/(x-9) y=(4+10)/(4-9) y=14/-5 y=-14/5 so the point is not containned on the equation...

OpenStudy (anonymous):

now with that sayd i think u can figure the rest out